OpenStudy (anonymous):

any thoughts on this one multiply and simplify the expression (3b-2)(4b-4)

OpenStudy (kainui):

Do FOIL method. Multiply First, Outer, Inner, and Last terms together and add them up.

OpenStudy (anonymous):

12b^2-12b-8b+8=12b^2-20b+8

jimthompson5910 (jim_thompson5910):

Or, you can do this (3b-2)(4b-4) 3b(4b-4) - 2(4b-4) I'll let you finish

OpenStudy (anonymous):

so is this even close 12b-18

jimthompson5910 (jim_thompson5910):

what is 3b(4b-4)

jimthompson5910 (jim_thompson5910):

what does it distribute out to

OpenStudy (anonymous):

12b-4

jimthompson5910 (jim_thompson5910):

3b times 4b = ???

OpenStudy (anonymous):

no wait 12b^2

jimthompson5910 (jim_thompson5910):

what is 3b times -4?

OpenStudy (anonymous):

-12

jimthompson5910 (jim_thompson5910):

-12b

jimthompson5910 (jim_thompson5910):

since that b gets tacked on the end

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

so 3b(4b-4) becomes 12b^2 - 12b

jimthompson5910 (jim_thompson5910):

do the same with the other group

OpenStudy (anonymous):

so now I am multiply 2 times 4b-4 correct

OpenStudy (anonymous):

so minues -8b-8 correct

OpenStudy (anonymous):

the final answer would look like this 12b^2-12b-8b-8 is this what the final answer should look like

jimthompson5910 (jim_thompson5910):

We want to distribute -2(4b - 4), so -2 times 4b = -8b -2 times -4 = 8 ---------------- This means -2(4b - 4) becomes -8b + 8

jimthompson5910 (jim_thompson5910):

So we now have (3b-2)(4b-4) 3b(4b-4) - 2(4b-4) 12b^2 - 12b - 8b + 8

jimthompson5910 (jim_thompson5910):

that's not the final answer, but one step away from it

OpenStudy (anonymous):

12b^2-20b+8 Final

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

thank you very much

jimthompson5910 (jim_thompson5910):

np